# thanks: Corresponding author

### v1

@@ I Introduction
The mass $m_{a}$ of the QCD axion DM is expected to have a narrow viable range if the associated PQ symmetry is broken after inflation in the early Universe (see, for example, Ref. [10]). Cosmological and lattice calculations suggest that post-inflationary scenarios prefer an axion mass in the tens to hundreds of $\mu\mathrm{eV}$ [11, 12, 13, 14, 15, 16] range, with a recent analysis suggesting $18040~{}\mu{\rm eV}<m_{a}<$ 180 $\mu$eV [17]. This mass range can be well covered by resonant cavity technique proposed below, in the photon frequency range 0.5-50 GHz, which will be the focus of our discussion.
The most sensitive experimental approach to detecting axions in the above mass window is the haloscope technique, which exploits axion–photon conversion in a resonant microwave cavity permeated by a strong magnetic field, as was suggested first by Sikivie [7]. However, haloscopes face a fundamental limitation: as the resonant frequency increases, the physical volume of the cavity—and hence the signal power—must decrease. For an axion signal at 25 GHz, a natural cavity volume is $\sim 1~{}\mathrm{cm}^{3}$, over $10^{5}\times$ smaller than the volumes used in the original ADMX configuration [18, 19, 20, 21, 22]. Covering the 0.5-50 GHz range at half the resolution matched to the axion coherence bandwidth, $\Delta\nu/\nu\sim 10^{-6}$, requires $\sim 10^{7}$ distinct frequency channels (the band width for axions is set by their small kinetic energy $\propto v^{2}$, where $v\sim 10^{-3}$ is the virial velocity in our galactic neighborhood). This scaling imposes a stringent limitation on scanning speed, primarily because conventional readout electronics contribute significant quantum or thermal noise, requiring long integration times per frequency channel. For example, assuming a typical $\sim 100$ s integration time per bin, covering the above frequency range would take $\sim 30$ years. By contrast, our proposed axion signal probing method, based on laser readout through a Fabry-Pérot interferometer and electro-optic modulation, effectively bypasses the limitations of traditional RF electronics. It enables sensitivity well beyond the quantum noise floor typically imposed by the detection system. The method proposed here can potentially speed up scanning rates by more than an order of magnitude, using available technology.
@@
Figure 1: The current status of the axion to two photon coupling in the mass range 1-300 $\mu$eV, corresponding to the frequency range 0.2-70 GHz (from Ref. [23]). CARAMEL aims to facilitate probing the frequency range of 0.5-50 GHz, or $\sim$ (2-200) $\mu$eV, with better than DFSZ sensitivity within the next five years. CARAMEL can potentially cover the preferred post-inflationary parameter space, corresponding to 40-180 $\mu$eV [17], at better than DFSZ sensitivity, using presently available technical capabilities.
The current status of axion searches is summarized by the plot in Fig.1, taken from Ref. [23]. As can be seen here, the window corresponding to the preferred post-inflationary PQ axion DM, $\sim 40-180$ $\mu$eV, within the KSVZ [24, 25] and DFSZ [8, 26] scenarios, remains largely unexplored. This hints at the inherent challenges associated with this regime of parameters. Progress in this range has been slow due to the convergence of several technical challenges: the cavity volume decreases with increasing frequency, conventional copper cavities suffer from degraded quality factors at high frequencies, and quantum-limited linear amplifiers constrain the readout sensitivity. Meanwhile, single-photon detectors, though promising, are not yet mature for robust operation under realistic laboratory conditions.
To overcome these barriers, multiple high-frequency haloscope concepts are under active development. These include the pizza-cavity design [27], which partitions the cavity into multiple sectors to increase frequency while preserving volume efficiency; photonic crystal [28, 29] and metamaterial cavities [30], which decouple frequency from physical size; and horn-array haloscopes [31], which leverage large metal surface areas to achieve broadband, volume-efficient searches at high frequencies. Additionally, pioneering experiments such as MADMAX [32, 33, 34], and DALI [35] explore dielectric haloscope techniques using layered dielectric interfaces to coherently boost the axion-induced signal power. ALPHA using the electron plasma resonance frequency of metalic rods [36], and CAST-CAPP [37] using a decommissioned dipole magnet in an efficient geometry. Superconducting cavities with quality factors comparable to or exceeding the expected axion quality factor, even in the presence of strong magnetic fields ($\sim 8\,\mathrm{T}$ and above), have been demonstrated using high-temperature superconducting (HTS) tapes [38, 39, 40]. These developments indicate that this technology is maturing and offer a scalable path toward detecting axions in the tens to hundreds of $\mu\mathrm{eV}$ mass range, as motivated by post-inflation QCD axion models.
@@
Our proposal combines electro-optic (EO) detection [41] and a “resonant probing” technique [1], see also [42, 43, 44]. These methods, previously suggested independently, are now unified to offer:
@@ II Detection Principle
where $g_{a\gamma\gamma}=(\alpha/\pi)(g_{\gamma}/f_{a})$ is the axion-photon coupling, $\alpha$ is the fine structure constant, and $g_{\gamma}$ takes values of $-0.97$ (KSVZ model [24, 25]) or $0.36$ (DFSZ model [8, 26]). $\rho_{a}$ is the local axion dark matter density, and $\nu_{a}=m_{a}c^{2}/h$ is the axion Compton frequency ($c=2.998\times 10^{8}$ m/s is the speed of light). In the above, $\langle\mathbf{B}_{e}^{2}\rangle$ is the squared average external magnetic field, $V$ is the cavity volume, $Q_{a}\approx 10^{6}$ is the axion quality factor, and $Q_{c}$ is the cavity unloaded quality factor. The form factor $G$ reflects the mode-dependent overlap of the axion-induced field with the resonant mode and is approximately 0.69 for the TM010 mode.
@@
$E_{a}=\sqrt{\frac{P_{a}Q_{c}}{\omega\epsilon_{0}V}}\,,$ (7)
@@
For an integration time of $t=3$ seconds, assumed henceforth, we get
@@
which would yield an $\text{SNR}_{\rm FP}\sim 0.04$, which is insufficient for detection. However, we will see next how using a probe radio-frequency (RF) beam at the axion frequency of oscillation can further enhance the signal.
@@ II.1 Probing Method with Electro-Optic Readout
Substituting the above benchmark values, we find $\psi_{\rm probe}\approx 7.7\times 10^{-11}$ rad. As before, if we employ a FP cavity with a finesse ${\cal F}$, we can enhance the signal
@@ II.2 RF probe effect on sensitivity
valid when $Q_{c}\leq Q_{a}$. A more general formula—valid across both classical and quantum limits and for arbitrary cavity quality factors—is derived in the Supplemental Material, Eq. 52
@@
This expression smoothly interpolates between the quantum regime at low temperatures (where $\coth\to 1$) and the classical limit at high temperatures (where $\coth\to\frac{2k_{B}T}{\hbar\omega}$).
@@ III Discussion: Surpassing the Laser Quantum Noise Limit
Until practical single-photon detectors become available in the GHz regime [50, 51, 52, 53, 50, 52, 54, 55, 56, 57], the method presented here offers a compelling enhancement in both sensitivity and scanning speed. Specifically, this approach can accelerate axion searches by more than an order of magnitude through two mechanisms. First, by injecting a coherent probe tone and detecting the axion-induced beat signal via an EO crystal, the effective SNR exceeds the noise limit of conventional haloscopes by a factor of 5–10 [58, 59, 60, 18, 61, 20, 62, 63, 64, 65, 66, 67, 68, 69, 37, 70, 71, 72, 73, 29]. Since the scanning rate scales inversely with the square of the SNR, this corresponds to a 25–100$\times$ increase in scanning speed while still achieving DFSZ sensitivity, even with the modest integration time of 3 seconds per frequency step assumed throughout this work. Second, the same Fabry–Pérot laser readout system can be used uniformly across a wide frequency range—from 0.5 GHz to at least 50 GHz—without reconfiguring the detection chain. This broadband capability is a critical advantage over conventional linear amplifiers and quantum-limited photon counters, which typically require frequency-specific optimization. Together, these features make the present method a versatile and powerful strategy for exploring the high-frequency axion parameter space that remains largely open.
@@
IV Summary
In summary, the CARAMEL approach unifies optical and RF techniques to achieve axion dark matter detection with enhanced sensitivity beyond the quantum noise limit. Its ability to operate across a broad frequency range, with low optical power and without complex quantum amplifiers, opens a scalable path forward in the high-frequency regime. The method is complementary to existing axion haloscopes and provides a versatile platform for future axion dark matter searches. The estimated reach of the method described in this work can cover much, and perhaps all [17], the relevant parameter space for post-inflationary axion dark matter, and potentially lead to a breakthrough in our understanding of puzzles in both the sub-atomic realm of hadronic interactions, and the large scale structure of the Universe.
@@ Acknowledgements.
V Supplemental Material
VI Signal-to-Noise Ratio Dependence on Quality Factor and Temperature
@@ 4. The Effect of Cavity Narrowness on SNR in the $Q_{c}\gg Q_{a}$ Regime
which can also result from Eq. 43 above for $Q_{c}\gg Q_{a}$. The integrated SNR is now proportional to $\sqrt{Q_{a}t/T}$ and independent of the cavity $Q_{c}$. The SNR scaling as $\sqrt{Q_{c}}$ holds only when $Q_{c}\lesssim Q_{a}$. For $Q_{c}\gg Q_{a}$, two competing effects emerge: while the cavity field amplitude continues to increase with $Q_{c}$, so does the thermal noise field, since the narrower cavity mode enhances thermal fluctuations per unit bandwidth despite admitting less total bandwidth. Simultaneously, only a fraction $Q_{a}/Q_{c}$ of the axion power spectrally overlaps with the cavity mode. These effects combine to cancel any further SNR gain, leading to saturation as $Q_{c}$ increases. The optimal regime is therefore $Q_{c}\approx Q_{a}$, where both the axion power and the thermal noise are spectrally matched. Future advances in single-photon detectors will be critical for pushing axion sensitivity further, particularly at low temperatures since they don’t suffer from limitations due to zero point fluctuations [50, 51, 52, 53, 50, 52, 54, 55, 56, 57].

### v2

@@ I Introduction
[…]
The most sensitive experimental approach to detecting axions in the above mass window is the haloscope technique, which exploits axion–photon conversion in a resonant microwave cavity permeated by a strong magnetic field, as was suggested first by Sikivie [7]. However, haloscopes face a fundamental limitation: as the resonant frequency increases, the physical volume of the cavity—and hence the signal power—must decrease. For an axion signal at 25 GHz, a natural cavity volume is $\sim 1~\mathrm{cm}^{3}$, over $10^{5}\times$ smaller than the volumes used in the original ADMX configuration [18, 19, 20, 21, 22, 23]. Covering the 0.5-50 GHz range at half the resolution matched to the axion coherence bandwidth, $\Delta\nu/\nu\sim 10^{-6}$, requires $\sim 10^{7}$ distinct frequency channels (the band width for axions is set by their small kinetic energy $\propto v^{2}$, where $v\sim 10^{-3}$ is the virial velocity in our galactic neighborhood). This scaling imposes a stringent limitation on scanning speed, primarily because conventional readout electronics contribute significant quantum or thermal noise, requiring long integration times per frequency channel. For example, assuming a typical $\sim 100$ s integration time per bin, covering the above frequency range would take $\sim 30$ years. By contrast, our proposed axion signal probing method, based on laser readout through a Fabry-Pérot interferometer and electro-optic modulation, effectively bypasses the limitations of traditional RF electronics. It enables sensitivity well beyond the quantum noise floor typically imposed by the detection system. The method proposed here can potentially speed up scanning rates by more than an order of magnitude, using available technology.
@@
[…]
To overcome these barriers, multiple high-frequency haloscope concepts are under active development. These include the pizza-cavity design [73], which partitions the cavity into multiple sectors to increase frequency while preserving volume efficiency; photonic crystal [74, 75] and metamaterial cavities [76], which decouple frequency from physical size; and horn-array haloscopes [77], which leverage large metal surface areas to achieve broadband, volume-efficient searches at high frequencies. Additionally, pioneering experiments such as MADMAX [78, 79, 80], and DALI [81] explore dielectric haloscope techniques using layered dielectric interfaces to coherently boost the axion-induced signal power. The ORGAN experiment [82, 83, 84] utilizes tunable resonant cavities in high magnetic fields to target axions in the 15–50 GHz range with ultra-low-noise quantum-limited readout systems. ALPHA uses the electron plasma resonance frequency of metalic rods [85], and CAST-CAPP [86] employed a decommissioned dipole magnet in an efficient geometry. Superconducting cavities with quality factors comparable to or exceeding the expected axion quality factor, even in the presence of strong magnetic fields ($\sim 8\,\mathrm{T}$ and above), have been demonstrated using high-temperature superconducting (HTS) tapes [87, 88, 89]. These developments indicate that this technology is maturing and offers a scalable path toward detecting axions in the tens to hundreds of $\mu\mathrm{eV}$ mass range, as motivated by post-inflation QCD axion models.
@@
Our proposal combines electro-optic (EO) detection [90] and a “resonant probing” technique [1]; see also [91, 92, 93]. These methods, previously suggested independently, are now unified to offer:
@@ II Detection Principle
where $g_{a\gamma\gamma}=(\alpha/\pi)(g_{\gamma}/f_{a})$ is the axion-photon coupling, $\alpha$ is the fine structure constant, and $g_{\gamma}$ takes values of 0.97-0.97 (KSVZ model [70, 71]) or $0.36$ (DFSZ model [8, 72]). $\rho_{a}$ is the local axion dark matter density, and $\nu_{a}=m_{a}c^{2}/h$ is the axion Compton frequency ($c=2.998\times 10^{8}$ m/s is the speed of light). In the above, $\langle\mathbf{B}_{e}^{2}\rangle$ is the squared average external magnetic field, $V$ is the cavity volume, $Q_{a}\approx 10^{6}$ is the axion quality factor, and $Q_{c}$ is the cavity unloaded quality factor. The form factor $G$ reflects the mode-dependent overlap of the axion-induced field with the resonant mode and is approximately 0.69 for the TM010 mode.
@@
$E_{a}=\sqrt{\frac{P_{a}Q_{c}}{\omega\epsilon_{0}V}}\,.$ (7)
@@
For an integration time of $t=3$ seconds, assumed henceforth, we get; see Appendix A:
@@
that would yield an $\text{SNR}_{\rm FP}\sim 0.04$, which is insufficient for detection. However, we will see next how using a probe radio-frequency (RF) beam at the axion frequency of oscillation can further enhance the signal.
@@ II.1 Probing Method with Electro-Optic Readout
[…]
@@ II.2 RF probe effect on sensitivity
valid when $Q_{c}\leq Q_{a}$. A more general formula—valid across both classical and quantum limits and for arbitrary cavity quality factors—is derived in Appendix B, (Eq. 59):
@@
This expression smoothly interpolates between the quantum regime at low temperatures [where $\coth(\hbar\omega/2k_{B}T)\to 1$ as $T\to 0$] and the classical limit at high temperatures [where $\coth(\hbar\omega/2k_{B}T)\to 2k_{B}T/\hbar\omega$ as $T\to\infty$].
@@ III Discussion: Surpassing the Laser Quantum Noise Limit
[…]
@@
IV Concluding Remarks and Summary
The CARAMEL method proposed in this work achieves a decisive advantage by replacing quantum-limited microwave amplification with a low-noise optical readout, even though the axion-photon conversion mechanism and the resonator design are conceptually identical to those used in traditional haloscope experiments. This readout scheme leverages the EO effect to convert weak electric field modulations into laser polarization signals, which can be measured with shot-noise-limited sensitivity over narrow detection bandwidths. In doing so, CARAMEL retains all the standard haloscope principles while extending their reach through a fundamentally different detection chain.
In addition to its low-noise readout advantage, CARAMEL offers two notable system-level benefits. First, it enables a unified detection platform across a vast frequency range, from sub-GHz to tens of GHz, without the need for reconfiguring quantum-limited amplifiers or microwave components. Second, it eliminates the added noise introduced by microwave amplification, which—even at the quantum limit—contributes at least half a photon of noise power to the system. The optical readout scheme in CARAMEL bypasses this fundamental amplifier constraint, allowing high-sensitivity measurements without incurring quantum-limited amplification penalties. We note that the CARAMEL approach does not beat vacuum noise in the absolute sense, but it resolves an axion signal in a narrow bin where vacuum noise is statistically small. This is achieved through the combined use of probing, Fabry–Pérot enhancement, and long integration. Thus, our method does not eliminate vacuum fluctuations — it resolves the axion field against them using time and coherence.
In summary, CARAMEL unifies optical and RF techniques to achieve axion dark matter detection with enhanced sensitivity beyond the quantum noise limit. Its ability to operate across a broad frequency range, with low optical power and without complex quantum amplifiers, opens a scalable path forward in the high-frequency regime. Our approach is complementary to existing axion haloscopes and provides a versatile platform for future axion dark matter searches. The method described in this work is projected to cover much – perhaps all [17] – the relevant parameter space for post-inflationary axion dark matter, and may lead to a breakthrough in resolving puzzles regarding both the sub-atomic interactions of hadrons and the large scale structure of the Universe.
@@ Acknowledgements.
Appendix A Laser Signal-to-Noise Ratio Estimation
We estimate the signal-to-noise ratio (SNR) for the axion-induced ellipticity in the probing method with optical readout.
The shot-noise-limited ellipticity sensitivity of the laser readout system is given by
$\delta\psi_{\text{shot}}=\sqrt{\frac{2h\nu}{P}}\approx 6.1\,\text{nrad}/\sqrt{\text{Hz}},$ (33)
where $\nu$ is the laser frequency, $P$ the laser power.
The axion-induced ellipticity is enhanced by the externally injected RF probing field and the finesse of the optical cavity. For the chosen parameters, the expected signal amplitude is
$\psi_{a}\approx 490\,\text{nrad}.$ (34)
The axion field has a finite coherence time of approximately
$\tau_{a}=\frac{1}{\Delta\nu_{a}}\approx 100\,\mu\text{s},$ (35)
corresponding to a bandwidth of $\Delta\nu_{a}=10\,\text{kHz}$. During a total integration time $t=3\,\text{s}$, the number of independent coherence segments is
$N=\frac{t}{\tau_{a}}=t\cdot\Delta\nu_{a}=3\,\text{s}\times 10^{4}\,\text{Hz}=3\times 10^{4}.$ (36)
The shot noise within each coherence time is
$\delta\psi_{\text{seg}}=\delta\psi_{\text{shot}}\cdot\sqrt{\Delta\nu_{a}}=6.1\,\text{nrad}\cdot 100=610\,\text{nrad}.$ (37)
Averaging incoherently over $N$ segments, the total noise decreases as
$\delta\psi=\frac{610\,\text{nrad}}{\sqrt{3\times 10^{4}}}\approx 3.5\,\text{nrad}.$ (38)
Therefore, the final SNR is given by
$\text{SNR}=\frac{\psi_{a}}{\delta\psi}=\frac{490\,\text{nrad}}{3.5\,\text{nrad}}\approx 140.$ (39)
This confirms that the probing method with optical readout can detect the axion-induced signal at high significance within a short integration time, provided the axion frequency is on resonance.
Appendix B Signal-to-Noise Ratio Dependence on Quality Factor and Temperature
@@ 4. The Effect of Cavity Narrowness on SNR in the $Q_{c}\gg Q_{a}$ Regime
which can also result from Eq. 50 above for $Q_{c}\gg Q_{a}$. The integrated SNR is now proportional to $\sqrt{Q_{a}t/T}$ and independent of the cavity $Q_{c}$. The SNR scaling as $\sqrt{Q_{c}}$ holds only when $Q_{c}\lesssim Q_{a}$. For $Q_{c}\gg Q_{a}$, two competing effects emerge: while the cavity field amplitude continues to increase with $Q_{c}$, so does the thermal noise field, since the narrower cavity mode enhances thermal fluctuations per unit bandwidth despite admitting less total bandwidth. Simultaneously, only a fraction $Q_{a}/Q_{c}$ of the axion power spectrally overlaps with the cavity mode. These effects combine to cancel any further SNR gain, leading to saturation as $Q_{c}$ increases. The optimal regime is therefore $Q_{c}\approx Q_{a}$, where both the axion power and the thermal noise are spectrally matched. Future advances in single-photon detectors will be critical for pushing axion sensitivity further, particularly at low temperatures since they do not suffer from limitations due to zero point fluctuations [98, 99, 100, 101, 98, 100, 102, 103, 104, 105].
@@ 5. General SNR Expression Including Quantum Fluctuations
Appendix C Bandwidth Considerations via RF Probing in EO Axion Detection
In the electro-optic (EO) probing method for axion dark matter detection, the injection of a coherent radiofrequency (RF) tone into the cavity radically alters the nature of the signal readout. Most importantly, it enables a dramatic reduction of the effective detection bandwidth, directly improving the signal-to-noise ratio (SNR) and enabling sensitivity beyond the conventional quantum limit associated with microwave amplifiers. However, the zero-point fluctuations of the electromagnetic field (vacuum noise) inside the cavity still remain.
1. Principle of RF Probing
In traditional haloscope experiments, the axion-induced signal appears as a weak oscillating electric field with a narrow intrinsic bandwidth determined by the velocity dispersion of the dark matter halo:
$\Delta\nu_{a}=\frac{\nu_{a}}{Q_{a}}\sim 10^{-6}\nu_{a}\sim 10\,\text{kHz},$ (61)
for axion frequencies $\sim 10$ GHz.
Without RF probing, one must integrate over this full axion bandwidth to recover the signal power, with noise accumulating over the same bandwidth.
By injecting a coherent RF tone $E_{p}\cos{(\omega_{p}t)}$ near the expected axion frequency $\omega_{a}$, the axion field $E_{a}\cos{(\omega_{a}t+\phi(t))}$ interferes with it to generate a beat signal at the difference frequency:
$\omega_{b}=|\omega_{a}-\omega_{p}|\sim\text{few kHz},$ (62)
This beat modulates the birefringence of the EO crystal, and appears as a low-frequency modulation in the polarization of a transmitted laser beam.
2. Detection Bandwidth from Integration Time
The probing technique concentrates the axion-induced signal into a narrow spectral line. The detection bandwidth is not set by $\Delta\nu_{a}$, but by the inverse of the integration time:
$\Delta\nu_{\text{det}}=\frac{1}{t}.$ (63)
For $t=3$ s, $\Delta\nu_{\rm det}\sim 0.33$  Hz. This allows one to measure a variance signal in a very narrow frequency bin, thereby reducing noise contribution.
3. Vacuum Noise Is Present and Probed
Vacuum fluctuations in the cavity, with RMS field:
$E_{\text{vac}}=\sqrt{\frac{\hbar\omega}{2\epsilon_{0}V}},$ (64)
are broadband and random-phase. When a coherent probe field is injected, vacuum noise beats with it just like the axion field does:
$E(t)=E_{\text{probe}}\cos(\omega t)+\delta E(t),$ (65)
with $\delta E(t)$ including vacuum, thermal, and axion components.
The ellipticity is proportional to:
$\psi(t)\propto E(t)^{2}\propto 2E_{\text{probe}}\delta E(t),$ (66)
which includes vacuum-induced and axion-induced modulation.
Therefore the EO method does not eliminate vacuum noise — it still contributes to the measured ellipticity fluctuations.
4. Why CARAMEL Can Still Resolve the Axion
What CARAMEL does is leverage:
  • Coherent RF probing to upconvert the axion signal into a narrow beat note;
  • Fabry–Pérot enhancement to increase signal ellipticity;
  • Narrowband optical readout $\Delta\nu_{\rm det}\ll\Delta\nu_{a}\leq\Delta\nu_{\rm cav}$ to suppress broadband noise.
Therefore, although the axion field is weaker than the vacuum field (e.g. $E_{a}/E_{\rm vac}\sim 10^{-2}-10^{-1}$) it is concentrated in a single narrow bin, while the vacuum noise is spread across $\Delta\nu\sim\nu/Q_{c}$. The effective vacuum noise in the detection bin is:
$E_{\text{vac,eff}}=E_{\text{vac}}\cdot\sqrt{\frac{\Delta f_{\text{det}}}{\Delta f_{\text{cavity}}}},$ (67)
which can be $\sim 10^{2}-10^{3}$ times smaller than the total RMS vacuum field.